OpenStudy (anonymous):

Evaluate the function f (x) = 4 • 7x for x = -1 and x = 2. Show your work.

OpenStudy (diamondboy):

k do u know how to plug in numbers?

OpenStudy (anonymous):

no

OpenStudy (trojanpoem):

f(-1) = 4 * 7 * -1 = - 28 can you do the x =2 ?

OpenStudy (diamondboy):

to plug in simply means to substitute

OpenStudy (diamondboy):

wherever u see x in d equation put the values of x u were given and simplify

OpenStudy (trojanpoem):

f(2) = 4* 7 * 2 = 56

OpenStudy (anonymous):

f(2)=4*7*2=56

OpenStudy (trojanpoem):

Right !

OpenStudy (anonymous):

thankx

OpenStudy (diamondboy):

I think trojan has it for u @jimmyjoew

OpenStudy (anonymous):

ok thankx for u too

OpenStudy (trojanpoem):

I didn't I showed him how to solve it and he solved the other himself.

OpenStudy (trojanpoem):

You are welcome.

OpenStudy (diamondboy):

yea that was what I meant